### v1

# TokaMark: A Comprehensive Benchmark for MAST Tokamak Plasma Models

Abstract.
Development and operation of commercially viable fusion energy reactors such as tokamaks require accurate predictions of plasma dynamics from sparse, noisy, and incomplete sensors readings. The complexity of the underlying physics and the heterogeneity of experimental data pose formidable challenges for conventional numerical methods, while simultaneously highlights the promise of modern data-native AI approaches. A major obstacle in realizing this potential is, however, the lack of curated, openly available datasets and standardized benchmarks. Existing fusion datasets are scarce, fragmented across institutions, facility-specific, and inconsistently annotated, which limits reproducibility and prevents a fair and scalable comparison of AI approaches. In this paper, we introduce TokaMark, a structured benchmark to evaluate AI models on real experimental data collected from the Mega Ampere Spherical Tokamak (MAST). TokaMark provides a comprehensive suite of tools designed to (i) unify access to multi-modal heterogeneous fusion data (ii) harmonize formats, metadata, temporal alignment and evaluation protocols to enable consistent cross-model and cross-task comparisons. The benchmark includes a curated list of 14 tasks spanning a range of physical mechanisms, exploiting a variety of diagnostics and covering multiple target use cases. A baseline model is provided to facilitate transparent comparison and validation within a unified framework. By establishing a unified benchmark for both the fusion and AI-for-science communities, TokaMark  aims to accelerate progress in data-driven plasma AI modeling, contributing to the broader goal of achieving sustainable and stable fusion energy. The benchmark, documentation, and tooling will be fully open sourced upon acceptance to encourage community adoption and contribution.
AI for science, Fusion benchmark, Data-driven plasma modeling, Multi-modal learning, MAST tokamak
copyright: nonecopyright: nonecopyright: noneccs: Hardware Renewable energyccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Open source modelccs: Information systems Extraction, transformation and loadingccs: Information systems Evaluation of retrieval results
1. Introduction
Nuclear fusion power is being explored as a potential long-term energy source with a unique combination of benefits. It offers the prospect of a carbon-neutral energy supply with abundant fuel and significant safety advantages over nuclear fission. However commercially viable fusion demands stable, sustained reactions that produce more power than the reactor consumes – a goal made difficult by the extreme physical conditions of thermonuclear confinement under which fusion must operate (Donné et al., 2025).
Fusion reactors such as tokamaks aim to reproduce on Earth the same process that powers the Sun, confining a plasma hotter than 100 million degrees Celsius using strong rapidly evolving magnetic fields because no material surface can survive direct contact with such plasma. This environment forces all measurements to be non-invasive and indirect, only partially inferring plasma state while the underlying dynamics evolve on microsecond-to-millisecond timescales. In this work, we dedicate our attention to the problem of modeling fusion plasma dynamics in tokamaks – one of the central challenges in fusion research. This problem encompasses a wide set of predictive tasks, including plasma shape and equilibrium inference, transport and profile evolution, and forecasting of MHD activity and disruptions.
1.1. AI for fusion plasma modeling
Traditional tokamak plasma modeling approaches rely either on first‑principles, physics‑driven solvers or on complex system identification pipelines. Although the governing physics is well understood, it is expressed through complex nonlinear partial differential equations whose key parameters are often unmeasured, only partially observable, or must be inferred indirectly from diagnostics. As a result, classical methods tend to be computationally intensive, brittle across operating regimes, and difficult to apply in real time. These challenges are further exacerbated by the nature of the experimental data itself, which reflects the extreme environment and sensing limitations of tokamak operation.
[…]
Tokamaks deploy a broad suite of heterogeneous diagnostics – magnetics, optical and X‑ray emission, microwave interferometry, and more – mounted on or behind the reactor walls and engineered to withstand heat, radiation, and electromagnetic stress. These sensors operate at widely different sampling rates, spatial resolutions, and noise characteristics producing data that is inherently heterogeneous, incomplete, multi‑rate and noisy.
Together, the complexity of the underlying physics and the heterogeneity of the data create an opportunity where modern AI methods can provide significant advantages. Data‑driven models are well suited to fusing multi‑modal diagnostics, learning latent plasma representations, and capturing nonlinear, multi‑scale temporal behavior. Unlike traditional solvers, AI models can operate directly on raw measurements, handle missing or asynchronous data, and produce accurate and efficient surrogates. This positions AI as a promising complement to physics‑based modeling for understanding, predicting, and ultimately controlling fusion plasma.
Previous work applying AI to fusion plasma problems have demonstrated promising results across variety of narrowly defined tasks, ranging from plasma shape reconstruction and profile forecasting to actuator optimization and disruptions prediction. However, most of those efforts use bespoke pipelines tailored to a small set of diagnostics, single device and single scientific objective, relying heavily on task-specific feature engineering and handcrafted labeling procedures. While successful within their respective scopes, these approaches typically optimize models for individual tasks and limit reuse across the experimental lifecycle. Much more work is needed to move beyond isolated point solutions toward broad, interoperable models capable of understanding fusion plasmas in a comprehensive way.
Inspired by the success of Foundation Models (FM) in language and vision, there is a growing expectation that analogous models trained on large corpora of tokamak data could learn rich, transferable plasma representations (Churchill, 2025). These models aim to internalize latent physical structure directly from data, serving as data‑driven analogs of first‑principles knowledge that can support a wide range of downstream tasks. Although still in early stages, this paradigm suggests a path toward generalist AI systems for fusion that complement physics-based modeling and reduce the need for handcrafted pipelines (Churchill, 2025).
Progress toward either specialized or generalist plasma models is hampered by the absence of open, standardized benchmarks. Fusion datasets remain fragmented across institutions, locked behind proprietary interfaces, or stored in domain‑specific formats that are difficult for Machine Learning researchers to access or interpret (Strand et al., 2022). Without unified task definitions, metrics, or evaluation protocols, it becomes impossible to provide fair comparison of methods and to measure progress systematically. A benchmark is therefore needed to frame the plasma modeling problems in a broader and more structured way, enabling cross‑comparison across algorithms, reproducibility across labs, and accessibility for researchers both inside and outside the fusion community.
At the same time, fusion diagnostic data provides unique set of challenge going beyond those found in standard datasets from more established areas of AI. The data is inherently multi‑modal, multi‑rate, and multi‑dimensional, with wide variations in signal quality, noise characteristics, and information content (see Figure 1 for an illustrative example). Many diagnostics operate asynchronously, contain gaps, or suffer partial failures; others produce high‑dimensional structured outputs such as flux maps or spectrograms. This complexity demands dedicated modeling approaches that can handle missing data, fuse heterogeneous signals, different modalities, and reason across fast and slow timescales – capabilities that are difficult to stress‑test without a standardized benchmark.
Taken together, these factors create a clear need for a comprehensive benchmark suite that defines common tasks, standardizes evaluation, and provides open access to representative fusion data. Establishing such a benchmark is essential for accelerating research, enabling fair comparison of models, and ultimately advancing data-driven plasma understanding and control.
1.2. TokaMark Benchmark Overview
To fill the need of comprehensive fusion plasma benchmarks we introduce TokaMark  the first large, open benchmark for evaluating AI models trained on real fusion diagnostics, built from recently released FAIR-MAST data (Jackson et al., 2025, 2024).
Data. To the best of our knowledge, FAIR-MAST is the only openly available dataset of real tokamak diagnostics. Recent releases (Jackson et al., 2025, 2024) curate a collection of real experiments and corresponding diagnostics measurements from MAST tokamak. From FAIR-MAST, we select 39 signals across heterogeneous modalities, harmonize metadata and build standardized loaders.
Tasks. For TokaMark we defined a diverse suite of 14 downstream tasks designed to probe core capabilities required of AI models for fusion plasmas: (i) representation learning from heterogeneous diagnostics; (ii) temporal reasoning across fast and slow timescales; (iii) robustness to incomplete state information; and (iv) generalization across operating regimes. Rather than optimizing for a single downstream objective, the tasks, organized into four groups, span a cascade of physical processes – from fast magnetic response to slower, transport-driven evolution and long-horizon precursors of MHD activity – while remaining closely aligned with routine experimental workflows. Wherever possible, tasks minimize reliance on expert-labeled targets, supporting self-supervised and weakly supervised formulations and enabling systematic evaluation of transferable plasma representations.
Evaluation. We introduce a hierarchical evaluation protocol aligned with the structure of FAIR-MAST and with scientific objectives. This hierarchy assesses both low-level prediction quality and high-level scientific meaning. For specialized models (single-task training), the hierarchy yields granular, signal-level diagnostics; for generalist models (FM-style pretraining across tasks), it provides modular comparisons across tasks and broad scientific objectives.
Baseline. Finally, we provide a strong yet accessible baseline: a multi-branch convolutional encoder–decoder architecture trained per each task that ingests heterogeneous inputs and predicts task-specific targets.
The following list summarizes our contributions:
  • Benchmark design. We define 14 tasks organized into four groups, with a hierarchical evaluation protocol, standardized windowing and error metrics.
  • Data packaging. We freeze a stable subset of FAIR-MAST, resolve schema inconsistencies, standardize metadata and units, and release it for reproducibility and long time compatibility.
  • Tools and API. We provide a Python package for task-specific data loading, processing and batching, masking and alignment utilities, and evaluation logic, integrated with PyTorch stack.
  • Baseline models. We release a multi-branch convolutional encoder–decoder baseline with reference configurations and training scripts, establishing reproducible baselines across all tasks (Seo et al., 2023) (Seo et al., 2024).
2. Preliminaries
To ground TokaMark benchmark in the reality of complex heterogeneous fusion data, we first introduce a signal taxonomy that organizes these measurements by function and modality, and allows us to provide a consistent vocabulary for describing inputs, targets, and tasks in the benchmark.
2.1. Data Taxonomy
Table 1. Signals taxonomy for TokaMark.
Category Subcategory/Signals Origin Frequency Modality
Magnetics Flux loops, pickup coils; saddle coils Sensor 5 kHz; 50 kHz Profile
Kinetics Thomson scattering; interferometer Sensor 0.2 kHz; 4 kHz Profile; time series
[…]
MHD Precursors Mirnov coils Sensor 500 kHz Profile
Currents Poloidal Field Coil currents; Solenoid current, Plasma current Sensor 4 kHz Profile; Time series
Voltages Poloidal Field Coil voltages Actuator 4 kHz Profile
References Reference Plasma Current, Reference Plasma Density Actuator 4 kHz Time series
Fueling NBI power, gas puffing Actuator 4 kHz Time series
Equilibrium Shape parameters, $\beta$ metrics; plasma boundary; flux map Derived 0.2 kHz Time series; profile; video
MAST is a spherically-shaped tokamak located in Culham, Oxfordshire (UK) and was operated by UKAEA and EURATOM from 1999 to 2013 (Sykes et al., 2001; Counsell et al., 2005; Meyer et al., 2009). MAST and tokamaks in general operate in short experimental cycles known as discharges or shots. The length of those cycles depends on the device size and in the case of MAST they are typically lasting around 2–3 seconds. Over its operational lifetime, MAST produced more than 30,000 shots, with each shot containing diagnostic signals measuring various properties of the plasma, including the magnetic field, plasma temperature, shape parameters, applied heating etc.
Recent works (Jackson et al., 2025, 2024) have created an open dataset of diagnostic data from a subset of the history of MAST. The FAIR-MAST dataset contains 11,573 shots from the last five experimental campaigns on MAST. In this work we utilized a total of 39 signals from FAIR-MAST in the design of our benchmark tasks.
We organize all signals along several complementary axes, summarized in Table 1. First, we group signals by category, which encodes their physical semantics–that is, the physical quantities they represent and how they are used by MAST workflows. This information corresponds to the first two columns of Table 1. Note that the second column contains both individual signals and subcategories. These intermediate groupings are introduced for brevity: they consolidate multiple signals that are typically used together (at least within the proposed benchmark) and share the same structural and functional properties. For example, the shape parameters subcategory includes several attributes jointly describing plasma geometry. A full list of the signals used in this work is provided in the appendix in Table 4.
Second, we distinguish signals by their origin. We define three classes: (i) sensors, corresponding to direct hardware measurements; (ii) actuators, representing controllable machine parameters used to steer plasma behavior; and (iii) derived signals, comprising quantities produced by reconstruction pipelines such as EFIT (e.g., shape parameters or flux maps).
The third axis is frequency, reflecting the wide range of sampling rates across diagnostics, and spanning from 0.2 kHz up to 500 kHz. Finally, the fourth axis is modality, describing the structural form of each signal: (i) time series, represented as 1D tensors (scalar value over time); (ii) profiles, represented as 2D tensors (vector over time); and (iii) videos, represented as 3D tensors (image over time). These latter two axes are particularly important for guiding the design of AI model architectures and loss functions.
2.2. Tasks Structural Taxonomy
The downstream tasks in TokaMark share a common structural formulation designed to reflect the online, window-based nature of plasma control and forecasting in tokamak experiments. Rather than operating on full-shot signals (as is often done in offline post-discharge analysis), our tasks are defined using input window and output window anchored at a reference time point.
Each task consists of one or more input signals and one or more output signals also referred to as targets. Inputs typically include both sensor signals (e.g., magnetics, radiatives, kinetics) and actuator signals (e.g., voltages, fueling), while outputs include either sensor signals or derived quantities like equilibrium reconstructions.
The alignment between the input and output windows naturally induces several families of modeling objectives:
  • Reconstruction. Given a set of sensors signals $A$ over the interval $t_{0}-\Delta_{input},t_{0}]$, the goal is to reconstruct a related set of signals $B$ over the same interval.
  • Autoregressive (AR) Forecasting. Given sensors signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ together with actuator trajectories over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, predict future values of the same signal set over $t_{0},t_{0}+\Delta_{output}]$.
  • Reconstructive (RC) Forecasting. Using sensors signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ and actuators over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, forecast a set of related outputs $B$ over $t_{0},t_{0}+\Delta_{output}]$ where $B$ may contain signals as in $A$.
We further distinguish tasks by their temporal dependency structure: Markovian tasks require only a short input window to make forecasts, reflecting fast dynamics, whereas Non-Markovian (NM) tasks require substantially longer input histories. The latter arise in phenomena governed by slower transport processes or evolving stability conditions, where even short-term predictions depend on extended plasma context.
Finally, TokaMark emphasizes self-supervised and weakly supervised task formulations whenever possible, reducing reliance on expert-labeled targets. This design was taken to encourage the development of AI models capable of discovering transferable plasma representations.
2.3. Data-driven Challenges
To complete the description of the FAIR-MAST dataset we highlight three core data-centric challenges that AI systems must overcome in order to resolve tasks in TokaMark. These challenges arise directly from the heterogeneous, multi-instrument nature of tokamak diagnostics and the realities of operating large fusion experiments.
  • Multi-fidelity. Diagnostic systems operate at widely varying sampling rates, ranging from a few hundred hertz to several hundred kilohertz. Resampling to a common high-frequency time base is computationally expensive and often unnecessary, while down-sampling to the lowest frequency discards information critical for resolving fast plasma phenomena. Effective models must therefore integrate and represent multi-rate temporal data without losing fidelity.
  • Multi-modality. Signals differ not only in physical meaning, units, numerical ranges, but also in structural forms of the corresponding data tensors. The dimensionality of the latter can be from 1 to 3 with the number of channel in each non-temporal dimension vary from 1 up to 170. AI systems must be able to fuse these diverse modalities into coherent representations suitable for downstream tasks.
  • Missing data. As with any experimental dataset, missing information is common. Entire signals may be absent for a given shot due to hardware issues. Individual signals may also contain missing time segments, for example due to limited acquisition windows or diagnostic failures. Naïvely discarding shots or windows with missing components wastes valuable examples and can introduce distributional bias. Robust approaches must therefore handle incomplete signals and irregular temporal coverage in a principled manner.
3. MAST Benchmark
In this section, we provide a description of the main components of TokaMark: tasks, data preparation and evaluation protocol.
3.1. Downstream Tasks
Table 2. Summary of groups and tasks for TokaMark.
Task Input Sensors Input Actuators Outputs Type Input Window Output Window
Group 1 1-1 Magnetics, Currents Shape parameters, $\beta$ metrics Reconstruction 5 ms 5 ms
1-2 Same as 1-1 Plasma boundary Reconstruction 5 ms 5 ms
1-3 Same as 1-1 Flux map Reconstruction 5 ms 5 ms
Group 2 2-1 Magnetics, Currents Voltages, NBI Power Currents, $\beta$ metrics, Shape parameters RC Forecasting 5 ms 25 ms
2-2 Same as 2-1 Same as 2-1 Plasma boundary RC Forecasting 5 ms 25 ms
2-3 Same as 2-1 Same as 2-1 Flux map RC Forecasting 5 ms 25 ms
Group 3 3-1 Thomson scattering References, Fueling Thomson scattering AR Forecasting 5 ms 50 ms
3-2 Thomson scattering, Radiatives Same as 3-1 Radiatives AR Forecasting 5 ms 50 ms
3-3 Magnetics, Currents, Interferometer Same as 3-1 Thomson scattering, $\beta$ metrics RC Forecasting, NM history 5 ms
Group 4 4-1 Magnetics, Mirnov coils, Currents, Radiatives, interferometer References, Fueling Soft X-ray AR Forecasting, NM history 100 ms
4-2 Magnetics, Mirnov coils, Currents, Radiatives, Kinetics Same as 4-1 Soft X-ray AR Forecasting, NM history 100 ms
4-3 Same as 4-1 Same as 4-1 Shape parameters RC Forecasting, NM history 100 ms
4-4 Same as 4-1 Same as 4-1 Plasma current AR Forecasting, NM history 100 ms
4-5 Same as 4-1 Same as 4-1 Mirnov diagnostics AR Forecasting, NM history 100 ms
For the TokaMark we constructed a suite of 14 tasks designed to evaluate the core capabilities expected of AI models for real-world fusion experiments: learning unified representations from heterogeneous diagnostics, temporal reasoning across both fast and slow dynamics, robustness to incomplete state information, and generalisation across operating regimes. Those tasks span across diverse timescales, modalities, and predictive objectives.
We arranged the tasks into 4 broad groups that represent some of the major modeling challenges an AI system must master: instantaneous reconstruction, short‑term magnetics dynamics, slow transport‑driven profile evolution, and long‑range forecasting of rare but safety‑critical events and MHD activity. Table 2 summarizes definitions of tasks and groups.
3.1.1. Group 1: Equilibrium Reconstruction
The tasks in group 1 define a suite of reconstruction problems where the objective is to infer plasma equilibrium – its shape, boundaries, and various properties – from instantaneous magnetic measurements. The targets of the three tasks in this group span multiple modalities from scalar parameters (task_1-1) to full plasma contour (task_1-2) and two-dimensional representations of the poloidal magnetic flux (task_1-3). The inputs of the tasks are identical and use standard magnetic diagnostics and actuator currents.
Equilibrium reconstruction is a foundational operation in tokamak experiments, performed routinely after every plasma discharge and forming the basis of essentially all downstream analysis. Traditionally, equilibrium is reconstructed by solving an inverse boundary-value problem governed by the Grad–Shafranov equation that describes the balance of magnetohydrodynamic force in axisymmetric tokamak plasmas. These solvers require multiple iterations for convergence and rely on parameterized assumptions about plasma profiles, making it challenging to deploy them in real time.
Group 1 tasks evaluate whether AI models can infer equilibrium structure directly from raw diagnostics, providing fast, numerics-free surrogates for a core analysis task that could underpin both offline interpretation and real-time control applications.
3.1.2. Group 2: Magnetics Dynamics
The tasks in Group 2 address short-timescale forecasting of magnetic signals, coil currents, and equilibrium evolution in response to applied actuator commands. Conceptually, these tasks parallel those in Group 1, but shift from static reconstruction to sequence-to-sequence forecasting with actuators signals now playing an essential role in the tasks inputs. Complexity increases in the three tasks in the group: from forecasting scalar quantities - such as plasma current evolution - to joint prediction of 2D equilibrium geometry over short horizons.
At these timescales, plasma evolution is dominated by inductive coupling between active coils, passive conducting structures, and the plasma current, together with the local response of the equilibrium to perturbations. These dynamics govern how the plasma reacts to control actions during a discharge and are central to plasma position, shape, and current control. While these effects are routinely modeled offline, a model capable of achieving real time latency would enable improved control policies.
Group 2 probes whether models can learn the effective plasma– machine response encoded in coupled circuit dynamics and plasma response. This would directly target capabilities required for closed-loop control, scenario planning, and digital twin applications.
3.1.3. Group 3: Profile Dynamics
Group 3 tasks focus on modeling the temporal evolution of kinetic profiles - primarily electron density and temperature - and on forecasting diagnostic signals associated with confinement-mode transitions. The three tasks in this group include short-horizon forecasting (task_3-1 and task_3-2) and pseudo-real-time reconstruction using partial, multi-rate diagnostic inputs (task_3-3).
Profile evolution is governed by transport physics acting on particle and energy balance, introducing slower characteristic timescales and intrinsic memory effects compared to magnetic dynamics. These processes control energy confinement and overall plasma performance and are central to understanding transport, optimising heating schemes, and predicting access to high-performance regimes. In practice, profile measurements are often sparse, delayed, or unavailable in real time.
The main goal of Group 3 tasks is to assess AI models ability to integrate incomplete diagnostic information over time to infer latent plasma state variables. This would support applications ranging from post-shot analysis, to integrated scenario design and real-time performance monitoring and control.
3.1.4. Group 4: MHD activity
Finally, the tasks in Group 4 comprise long-horizon forecasting of thermal quenches (task_4-1 and task_4-2), vertical displacement events (task_4-3), current quenches (task_4-4) and MHD activity leading to Locked Modes (task_4-5). These tasks are focused on early detection of plasma instability and disruptions precursors that emerge across multiple diagnostic modalities. These tasks require processing of rich, high frequency inputs–including magnetics, radiative diagnostics, and time–frequency MHD signatures—over extended temporal windows, exceeding well beyond immediate system response times.
The onset of MHD instabilities and disruptions is tightly linked to the evolving equilibrium and current distribution, even when the nonlinear dynamics of the instability itself are difficult to model explicitly. Disruption avoidance and mitigation are critical operational requirements for present-day tokamaks and future reactors, directly affecting machine protection. In practice, early-warning systems rely on detecting subtle precursors distributed across multiple diagnostics and evolving over extended time windows.
Group 4 tasks probe models abilities to integrate long-range temporal context and multi-modal information to anticipate loss-of-control events, testing capabilities that are essential for safe and reliable fusion operation.
3.2. Data Preparation
From the FAIR-MAST dataset we take all 11,573 available shots and extract the data for 39 signals required by out task definitions. Before being used for model training and evaluation, these signals undergo through a standardized preprocessing pipeline to ensure consistent formatting and alignment.
Data split. The benchmark dataset is divided into disjoint training, validation, and test subsets. To avoid information leakage across sets, the split is performed at the shot level representing independent experiments and employs random sampling technique. The split ratio is 80% / 10% / 10%, and all hyperparameter tuning is conducted exclusively on the training and validation sets. The test set is held out and used only for the final performance evaluation.
Window segmentation. Shot-level signals data is segmented into multiple input and output windows. The lengths of those windows are task dependent and specified in Table 2. Windows are extracted using a sliding-window approach with a stride of $0.001$ seconds between consecutive windows. Each input window is paired with its corresponding output window and treated as an independent window-level sample. We note that signals are used at their original sampling rates, and no resampling or imputation is performed.
Handling missing information. We impose additional constraints on the test set to ensure meaningful evaluation. Windows in which all input signals are entirely absent or composed solely of NaN values are excluded. Moreover, all output signals must be fully available over the entire prediction horizon.
3.3. Benchmark Evaluation
As part of the TokaMark, we introduce a hierarchical evaluation protocol for assessing model performance in a manner that aligns with scientific objectives and the structure of FAIR-MAST data. The design explicitly separates three levels of the hierarchy: (i) signals (individual physical quantities), and (ii) tasks (well-defined scientific goals) , and (iii) groups (broader physical objectives). This ensures that the benchmark assess both low-level prediction quality and high-level scientific utility. For low-level evaluation, the hierarchy provides granular signal-level insights that help diagnose which physics regimes the model captures or fails to capture. For models trained jointly across multiple tasks, the same hierarchy provides modular evaluations that highlight generality and transfer properties across tasks.
To respect the hierarchy, the evaluation aggregates errors according to the following progression:
$\text{samples}\;\rightarrow\;\text{windows}\;\rightarrow\;\text{signals}\;\rightarrow\;\text{tasks}\;\rightarrow\;\text{shots}.$
Samples are the atomic level of data. For a given task and shot, each data sample is denoted as $y_{k,i,j}$ and corresponds to a particular sample $j$ of the flattened data from window $i$ and signal $k$. In the following, $y_{k,i,j}$ denotes the ground truth value and $\hat{y}_{k,i,j}$ the corresponding model prediction.
Windows are containers of the equal per signal size storing $N_{k}$ samples. We compute a window-level Root-Mean-Square error error as:
(1) $\mathrm{RMSE}_{k,i}=\sqrt{\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}\left(y_{k,i,j}-\hat{y}_{k,i,j}\right)^{2}}$
Signals serve as first level of the evaluation hierarchy. For each signal $k$ containing $M_{k}$ windows, we aggregate all window errors via the signal-level Root-Mean-Square error and normalize it with the empirical standard deviation $\sigma_{k}$ computed for signal $k$ across all evaluation shots:
(2) $\mathrm{RMSE}_{k}=\sqrt{\frac{1}{M_{k}}\sum_{i=1}^{M_{k}}\mathrm{RMSE}_{k,i}^{2}},\quad\mathrm{NRMSE}_{k}=\frac{\mathrm{RMSE}_{k}}{\sigma_{k}}$
This yields a dimensionless quantity that expresses prediction error relative to the natural variability of the target making the metric comparable across signals. An error of $\mathrm{NRMSE}_{k}=1$ corresponds to a model no better than approximating the signal by its mean, while $\mathrm{NRMSE}_{k}<1$ indicates predictive value. We also note that $\mathrm{RMSE}_{k}$ and $\mathrm{NRMSE}_{k}$ are computed for a given task and shot.
Tasks. Task-level errors $\tilde{e}_{t}$ combine $K_{t}$ normalized output signal errors into a single score using a uniform mean, allowing one to quantify model performance on the scientific objective as a whole:
(3) $\tilde{e}_{t}=\frac{1}{K_{t}}\sum_{k=1}^{K_{t}}\mathrm{NRMSE}_{k}.$
Shots are the final step of the aggregation sequence. They represent distinct plasma discharges and therefore provide natural boundaries for independent experimental realizations. We report per-signal and per-task shot-level averaged scores correspondingly defined as
(4) $\mathrm{SignalScore}=\frac{1}{S}\sum_{s=1}^{S}\mathrm{NRMSE}_{s,k},\quad\mathrm{TaskScore}=\frac{1}{S}\sum_{s=1}^{S}\tilde{e}_{s,t}$
Here, $S$ is the total number of shots in the test set used for evaluation. Finally, the group level score is taken as an average of the corresponding task scores.
4. Baseline Model Experiments
The baseline model is a multi-branch convolutional architecture designed to handle heterogeneous spatio-temporal inputs and outputs with varying dimensionalities (see Figure 2). Each input modality is processed by a dedicated encoder (Seo et al., 2023), and each target variable is generated by a corresponding decoder, all connected through a shared latent representation.
4.1. Model Description
Encoders. For each input variable $v_{i}\in(L_{t}^{(0,v_{i})}\times L_{h}^{(0,v_{i})}\times L_{w}^{(0,v_{i})})$, an independent convolutional encoder is instantiated according to its modality: 1D convolutions for time series, 2D convolutions for profiles, and 3D convolutions for videos signals. Each encoder consists of a stack of $N$ convolutional layers with kernel size $K$, stride $s$, and padding $p$, followed by ReLU activations, max-pooling, and batch normalization. The feature maps are then flattened into a latent vector. The size of the feature map $L_{d}^{(\ell,v_{i})}$ along dimension $d\in(t,h,w)$ after encoder layer $\ell\in[1,N]$ and the flattened latent representation size $L^{(v_{i})}$ for variable $v_{i}$ can be computed as follows:
(5) $L_{d}^{(\ell,v_{i})}$ $=\left\lfloor\frac{1}{2}\left\lfloor\frac{L_{d}^{(\ell-1,v_{i})}+2p-K}{s}\right\rfloor+p-\frac{1}{2}\right\rfloor+1$
(6) $L^{(v_{i})}$ $=2^{N-1}DL_{t}^{(N,v_{i})}L_{h}^{(N,v_{i})}L_{w}^{(N,v_{i})}$
Figure 2. Multi-branch convolutional encoder–decoder
Latent fusion backbone. The latent embeddings produced by all encoder branches are concatenated and passed through a shared backbone. This backbone consists of a sequence of linear layers each followed by ReLU activation and dropout, progressively reducing the dimensionality back to $D$. This shared representation provides a compact summary of all input data and connects the encoders and decoders blocks.
Decoders. For each output variable $v_{o}\in(L_{t}^{(N,v_{o})}\times L_{h}^{(N,v_{o})}\times L_{w}^{(N,v_{o})})$, the decoder branch reconstructs the target output from the latent vector, mirroring the encoder structure. The shared latent vector is first reshaped into a compressed feature map of dimensionality $L^{(v_{o})}$, which is then progressively upsampled using transposed convolutions with the output padding $o_{p}$. Cropping operations are applied after each transposed convolution to ensure that the reconstructed outputs exactly match the required target dimensions $L_{d}^{(\ell,v_{o})}$. The sizes $L_{d}^{(\ell,v_{o})}$ of the feature map before decoder layer $\ell\in[0,N-1]$ and $L^{(v_{o})}$ of the compressed flattened feature map are:
(7) $L_{d}^{(\ell,v_{o})}$ $=\left\lceil s\cdot(L_{d}^{(\ell+1,v_{o})}-1)-2p+K+o_{p}\right\rceil$
(8) $L^{(v_{o})}$ $=2^{N-1}DL_{t}^{(0,v_{o})}L_{h}^{(0,v_{o})}L_{w}^{(0,v_{o})}$
4.2. Experimental Settings
Parameter Settings. For our architecture, we adopt a latent embedding dimension of $D=16$, with $N=3$ convolutional layers per encoder and decoder blocks. Each convolution uses a kernel size $K=3$, stride $s=3$, padding $p=1$ and output padding $o_{p}=1$. The same architectural hyperparameters — number of layers, kernel size, stride, and padding — are applied consistently across all tasks, establishing a unified baseline for the benchmark.
Data Preprocessing. We run a set of model specific preprocessing steps: first, we replace NaN values with zeros, second, we standardize data using signal-computed zero mean and unit variance scaler. This ensures numerical stability and helps to improve convergence during training. The training and validation data is down-sampled by taking window samples with a stride of $0.005$ seconds for Markovian tasks and $0.025$ seconds for non-Markovian tasks. Furthermore, because our architecture requires a fixed-length context, all input windows are truncated to a duration of 50 ms for non-Markovian tasks.
Training Procedure. We train our model on mini-batches containing data from 32 shots using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and early stopping with a patience of 10 epochs which terminates training if the validation loss does not improve over this period to prevent overfitting. We also employ a multi-output mean squared error loss, which averages the loss across all outputs of the model. Table 3 summarizes the size of models per task, computed from the encoder and decoder feature maps.
4.3. Experimental Results
The results of the baseline models evaluation on TokaMark tasks are presented in Table 3 containing task level scores and group averages. Those results demonstrate a clear distinction between tasks complexities: the model performs well for equilibrium reconstruction (Group 1) and magnetics dynamics (Group 2) tasks with group-level scores below 0.17 but degrades for profile dynamics (Group 3) and MHD activites (Group 4) tasks confirming that, as expected, these are more difficult tasks. Notably, even within the same group tasks scores vary substantially. For instance, plasma boundary reconstruction and forecasting tasks 1-2 and 2-2 are resolved much better then their counterparts within the respective groups. On the other side, the error score for task 4-5 is by far the worst and exceeds unity, suggesting the corresponding signals are poorly constrained or inadequately represented.
These results should be interpreted with a consideration to the nature of a baseline model: the architecture is generic, without physics-informed priors or task-specific tuning. The scores reported – especially those for profile dynamics and MHD activities tasks – highlight intrinsic benchmark difficulty rather than flaws in optimization. This baseline establishes a realistic lower bound and identifies areas for improvement, such as physics-aware modeling, task-adaptive weighting, or specialized architectures that should be considered by the users of TokaMark.
Table 3. TaskScore and GroupScore for TokaMark.
Task TaskScore Model size (M)
Task 1-1 0.1882 0.343
Task 1-2 0.0482 0.239
Task 1-3 0.2552 0.257
Group 1 Score 0.1656
Task 2-1 0.1551 0.436
Task 2-2 0.0517 0.277
Task 2-3 0.1724 0.295
Group 2 Score 0.1290
Task 3-1 0.3252 0.158
Task 3-2 0.3309 0.761
Task 3-3 0.3607 0.336
Group 3 Score 0.3380
Task 4-1 0.3445 1.518
Task 4-2 0.3311 1.571
Task 4-3 0.2702 0.672
Task 4-4 0.4292 0.687
Task 4-5 1.0053 4.741
Group 4 Score 0.4217
5. Conclusions
In this work, we introduced TokaMark, the first large-scale, open benchmark specifically designed for evaluating AI models on MAST tokamak diagnostics. We provided a complete open-source training and evaluation stack for 14 diverse downstream tasks together with a strong baseline model, creating an integrated framework for benchmarking, tooling, and model development.
TokaMark opens the door to more systematic and reproducible research in fusion plasma modeling. It provides a platform for exploring advanced representation learning of plasma, short- and long-horizon predictions, and generalization across tokamak operating regimes. We believe the adoption of TokaMark will accelerate progress toward practical, data-driven fusion models, foster stronger collaboration between the fusion and machine learning communities, and ultimately contribute to the development of stable and commercially viable fusion energy.

### v2

@@
 Development and operation of commercially viable fusion energy reactors such as tokamaks require accurate predictions of plasma dynamics from sparse, noisy, and incomplete sensors readings.
-The complexity of the underlying physics and the heterogeneity of experimental data pose formidable challenges for conventional numerical methods, while simultaneously highlights the promise of modern data-native AI approaches.
+The complexity of the underlying physics and the heterogeneity of experimental data pose formidable challenges for conventional numerical methods, while simultaneously highlight the promise of modern data-native AI approaches.
 A major obstacle in realizing this potential is, however, the lack of curated, openly available datasets and standardized benchmarks.
@@
 In this paper, we introduce TokaMark, a structured benchmark to evaluate AI models on real experimental data collected from the Mega Ampere Spherical Tokamak (MAST).
-TokaMark provides a comprehensive suite of tools designed to (i) unify access to multi-modal heterogeneous fusion data (ii) harmonize formats, metadata, temporal alignment and evaluation protocols to enable consistent cross-model and cross-task comparisons.
-The benchmark includes a curated list of 14 tasks spanning a range of physical mechanisms, exploiting a variety of diagnostics and covering multiple target use cases.
+TokaMark provides a comprehensive suite of tools designed to (i) unify access to multi-modal heterogeneous fusion data, and (ii) harmonize formats, metadata, temporal alignment and evaluation protocols to enable consistent cross-model and cross-task comparisons.
+The benchmark includes a curated list of 14 tasks spanning a range of physical mechanisms, exploiting a variety of diagnostics and covering multiple operational use cases.
 A baseline model is provided to facilitate transparent comparison and validation within a unified framework.
-By establishing a unified benchmark for both the fusion and AI-for-science communities, TokaMark  aims to accelerate progress in data-driven plasma AI modeling, contributing to the broader goal of achieving sustainable and stable fusion energy.
+By establishing a unified benchmark for both the fusion and AI-for-science communities, TokaMark  aims to accelerate progress in data-driven AI-based plasma modeling, contributing to the broader goal of achieving sustainable and stable fusion energy.
 The benchmark, documentation, and tooling will be fully open sourced upon acceptance to encourage community adoption and contribution.
-AI for science, Fusion benchmark, Data-driven plasma modeling, Multi-modal learning, MAST tokamak
-copyright: nonecopyright: nonecopyright: noneccs: Hardware Renewable energyccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Open source modelccs: Information systems Extraction, transformation and loadingccs: Information systems Evaluation of retrieval results
+AI for science, Data-driven modeling, Fusion benchmark, Plasma dynamics, Multi-modal learning, MAST tokamak
+copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 32nd SIGKDD Conference on Knowledge Discovery and Data Mining, 2026; August 09–13, 2026; Jeju, Koreaisbn: XXX-X-XXXX-XXXX-X/2026/08copyright: nonecopyright: nonecopyright: noneccs: Applied computing Physical sciences and engineeringccs: Computing methodologies Knowledge representation and reasoningccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Open source modelccs: Information systems Extraction, transformation and loading
 1. Introduction
@@
 It offers the prospect of a carbon-neutral energy supply with abundant fuel and significant safety advantages over nuclear fission.
-However commercially viable fusion demands stable, sustained reactions that produce more power than the reactor consumes – a goal made difficult by the extreme physical conditions of thermonuclear confinement under which fusion must operate (Donné et al., 2025).
-Fusion reactors such as tokamaks aim to reproduce on Earth the same process that powers the Sun, confining a plasma hotter than 100 million degrees Celsius using strong rapidly evolving magnetic fields because no material surface can survive direct contact with such plasma.
+However, commercially viable fusion demands stable, sustained operation that produces more energy than the power plant consumes—a goal made difficult by the extreme physical conditions of thermonuclear confinement under which fusion must operate (Donné et al., 2025).
+Magnetic confinement fusion reactors such as tokamaks aim to reproduce the process that powers the Sun in a controlled environment, confining a plasma hotter than 100 million degrees Celsius through the use of strong magnetic fields because no material surface can survive direct contact with such plasma (Wesson, 1987; Freidberg, 2007).
 This environment forces all measurements to be non-invasive and indirect, only partially inferring plasma state while the underlying dynamics evolve on microsecond-to-millisecond timescales.
-In this work, we dedicate our attention to the problem of modeling fusion plasma dynamics in tokamaks – one of the central challenges in fusion research.
-This problem encompasses a wide set of predictive tasks, including plasma shape and equilibrium inference, transport and profile evolution, and forecasting of MHD activity and disruptions.
+In this work, we dedicate our attention to the problem of modeling fusion plasma dynamics in tokamaks—one of the central challenges in fusion research.
+This problem encompasses a wide set of predictive tasks, including plasma shape and equilibrium inference, transport and profile evolution, and forecasting of magnetohydrodynamics (MHD) activity and disruptions.
 1.1. AI for fusion plasma modeling
-Traditional tokamak plasma modeling approaches rely either on first‑principles, physics‑driven solvers or on complex system identification pipelines.
-Although the governing physics is well understood, it is expressed through complex nonlinear partial differential equations whose key parameters are often unmeasured, only partially observable, or must be inferred indirectly from diagnostics.
-As a result, classical methods tend to be computationally intensive, brittle across operating regimes, and difficult to apply in real time.
-These challenges are further exacerbated by the nature of the experimental data itself, which reflects the extreme environment and sensing limitations of tokamak operation.
 […]
-Tokamaks deploy a broad suite of heterogeneous diagnostics – magnetics, optical and X‑ray emission, microwave interferometry, and more – mounted on or behind the reactor walls and engineered to withstand heat, radiation, and electromagnetic stress.
-These sensors operate at widely different sampling rates, spatial resolutions, and noise characteristics producing data that is inherently heterogeneous, incomplete, multi‑rate and noisy.
-Together, the complexity of the underlying physics and the heterogeneity of the data create an opportunity where modern AI methods can provide significant advantages.
-Data‑driven models are well suited to fusing multi‑modal diagnostics, learning latent plasma representations, and capturing nonlinear, multi‑scale temporal behavior.
+Traditional approaches to tokamak plasma modeling are rooted in well-established first-principles descriptions of magnetized plasma dynamics.
+These descriptions are expressed through coupled, nonlinear systems of partial differential equations, whose numerical solutions often requires high-fidelity, multi-scale simulations.
+While such models are indispensable for predictive studies, their computational cost severely limits their routine use, both in exploring the full phenomenology of plasma behavior, and in systematically interrogating large experimental datasets.
+In particular, many physically relevant regimes remain difficult to characterize in detail because comprehensive parameter scans and high-fidelity simulations are prohibitively expensive.
+The same computational burden complicates data-driven inference.
+Key parameters governing plasma behavior—such as transport coefficients, source terms, or stability-relevant profile features—are frequently unmeasured or only indirectly observable, and inferring them typically requires repeated forward simulations embedded within optimization or system identification loops.
+As a result, fitting models to experimental data becomes costly and brittle across operating regimes.
+These limitations make the direct use of first-principles solvers infeasible for real-time applications, where stringent latency constraints rule out iterative or high-fidelity numerical solutions.
+The challenge is further compounded by the experimental characteristics of tokamak data.
+Tokamaks deploy a broad suite of heterogeneous diagnostics–magnetics, optical and X‑ray emission, microwave interferometry, and more (Morris and the MAST team., 2002)—mounted on or behind the reactor walls and engineered to withstand extreme heat, radiation, and electromagnetic stress.
+These sensors operate at widely different sampling rates, spatial resolutions, and noise characteristics, producing data that is inherently heterogeneous, incomplete, multi‑rate, and noisy.
+Together, the complexity of the underlying physics and the heterogeneity of the data create an opportunity where modern AI methods can provide significant advantages and complement physics‑based modeling.
+Data‑driven models are well suited to fusing multi‑modal diagnostics, learning latent plasma representations, and capturing nonlinear, multi‑scale temporal behavior without explicit physical parameterization.
 Unlike traditional solvers, AI models can operate directly on raw measurements, handle missing or asynchronous data, and produce accurate and efficient surrogates.
-This positions AI as a promising complement to physics‑based modeling for understanding, predicting, and ultimately controlling fusion plasma.
-Previous work applying AI to fusion plasma problems have demonstrated promising results across variety of narrowly defined tasks, ranging from plasma shape reconstruction and profile forecasting to actuator optimization and disruptions prediction.
+However, these advantages come with caveats: learned models may fail silently when operating outside their training distribution, lack guaranteed physical consistency, and are generally more difficult to interpret.
+Previous work applying AI to fusion plasma problems have demonstrated promising results across a variety of narrowly defined tasks, ranging from plasma shape reconstruction (Wan et al., 2023; Wai et al., 2022; Rossi et al., 2023) and profile forecasting (Wan et al., 2021; Abbate et al., 2021, 2023, 2025; Char et al., 2024; Kit et al., 2024; Wakatsuki et al., 2023) to actuator optimization (Wang et al., 2025, 2024; Yang et al., 2020; Schramm et al., 2024; Seo et al., 2024, 2021, 2022; Vega et al., 2022; Degrave et al., 2022; Abbate et al., 2023; Orozco et al., 2022) and disruptions prediction (Zhang et al., 2020; Zhu et al., 2023; Priyanka et al., 2024; Rea et al., 2018; Lucas et al., 2024; Montes, 2021; Montes et al., 2021; Churchill et al., 2020; Ferreira et al., 2020; Guo et al., 2021; Kates-Harbeck et al., 2019; Zhu et al., 2021b; De Vries et al., 2011; Zhu et al., 2021a; Aymerich et al., 2022).
 However, most of those efforts use bespoke pipelines tailored to a small set of diagnostics, single device and single scientific objective, relying heavily on task-specific feature engineering and handcrafted labeling procedures.
-While successful within their respective scopes, these approaches typically optimize models for individual tasks and limit reuse across the experimental lifecycle.
-Much more work is needed to move beyond isolated point solutions toward broad, interoperable models capable of understanding fusion plasmas in a comprehensive way.
+While successful within their respective scopes, these approaches typically optimize models for individual tasks, which limits reuse across the experimental lifecycle.
+Adding to ongoing efforts (Dong et al., 2025; Yang et al., 2025), more work is needed to move beyond isolated point solutions toward broad, interoperable models capable of understanding fusion plasmas in a comprehensive way.
 Inspired by the success of Foundation Models (FM) in language and vision, there is a growing expectation that analogous models trained on large corpora of tokamak data could learn rich, transferable plasma representations (Churchill, 2025).
@@
 Fusion datasets remain fragmented across institutions, locked behind proprietary interfaces, or stored in domain‑specific formats that are difficult for Machine Learning researchers to access or interpret (Strand et al., 2022).
-Without unified task definitions, metrics, or evaluation protocols, it becomes impossible to provide fair comparison of methods and to measure progress systematically.
+Without unified task definitions, metrics, or evaluation protocols, it becomes impossible to provide a fair comparison of methods and to measure progress systematically.
 A benchmark is therefore needed to frame the plasma modeling problems in a broader and more structured way, enabling cross‑comparison across algorithms, reproducibility across labs, and accessibility for researchers both inside and outside the fusion community.
-At the same time, fusion diagnostic data provides unique set of challenge going beyond those found in standard datasets from more established areas of AI.
+At the same time, fusion diagnostic data provide a unique set of challenges going beyond those found in standard datasets from more established areas of AI.
 The data is inherently multi‑modal, multi‑rate, and multi‑dimensional, with wide variations in signal quality, noise characteristics, and information content (see Figure 1 for an illustrative example).
 Many diagnostics operate asynchronously, contain gaps, or suffer partial failures; others produce high‑dimensional structured outputs such as flux maps or spectrograms.
-This complexity demands dedicated modeling approaches that can handle missing data, fuse heterogeneous signals, different modalities, and reason across fast and slow timescales – capabilities that are difficult to stress‑test without a standardized benchmark.
+This complexity demands dedicated modeling approaches that can handle missing data, fuse heterogeneous signals, different modalities, and reason across fast and slow timescales—capabilities that are difficult to stress‑test without a standardized benchmark.
 Taken together, these factors create a clear need for a comprehensive benchmark suite that defines common tasks, standardizes evaluation, and provides open access to representative fusion data.
@@
 1.2. TokaMark Benchmark Overview
-To fill the need of comprehensive fusion plasma benchmarks we introduce TokaMark  the first large, open benchmark for evaluating AI models trained on real fusion diagnostics, built from recently released FAIR-MAST data (Jackson et al., 2025, 2024).
+To fill the need for comprehensive fusion plasma benchmarks, we introduce TokaMark, the first large, open benchmark for evaluating AI models trained on real fusion data.
 Data.
-To the best of our knowledge, FAIR-MAST is the only openly available dataset of real tokamak diagnostics.
+To the best of our knowledge, FAIR-MAST data represents the only openly available dataset of real tokamak diagnostics.
 Recent releases (Jackson et al., 2025, 2024) curate a collection of real experiments and corresponding diagnostics measurements from MAST tokamak.
@@
 Tasks.
-For TokaMark we defined a diverse suite of 14 downstream tasks designed to probe core capabilities required of AI models for fusion plasmas: (i) representation learning from heterogeneous diagnostics; (ii) temporal reasoning across fast and slow timescales; (iii) robustness to incomplete state information; and (iv) generalization across operating regimes.
-Rather than optimizing for a single downstream objective, the tasks, organized into four groups, span a cascade of physical processes – from fast magnetic response to slower, transport-driven evolution and long-horizon precursors of MHD activity – while remaining closely aligned with routine experimental workflows.
-Wherever possible, tasks minimize reliance on expert-labeled targets, supporting self-supervised and weakly supervised formulations and enabling systematic evaluation of transferable plasma representations.
+For TokaMark, we defined a diverse suite of 14 downstream tasks organized into 4 groups, designed to probe core capabilities required by AI models for fusion plasmas: (i) representation learning from heterogeneous diagnostics; (ii) temporal reasoning across fast and slow timescales; (iii) robustness to incomplete state information; and (iv) generalization across operating regimes.
+Rather than optimizing for a single downstream objective, the tasks span a cascade of physical processes—from fast magnetic response to slower, transport-driven evolution and long-horizon precursors of MHD activity—while remaining closely aligned with routine experimental workflows.
+Wherever possible, the tasks minimize reliance on expert-labeled targets, supporting self-supervised and weakly supervised formulations, and enabling systematic evaluation of transferable plasma representations.
 Evaluation.
 We introduce a hierarchical evaluation protocol aligned with the structure of FAIR-MAST and with scientific objectives.
-This hierarchy assesses both low-level prediction quality and high-level scientific meaning.
-For specialized models (single-task training), the hierarchy yields granular, signal-level diagnostics; for generalist models (FM-style pretraining across tasks), it provides modular comparisons across tasks and broad scientific objectives.
+This hierarchy assesses both low-level prediction quality and high-level scientific utility.
+For specialized models (single-task training), the hierarchy yields granular, signal-level diagnostics; for generalist models (FM-style pretrained and then fine-tuned across tasks), it provides modular comparisons across tasks and broad scientific objectives.
 Baseline.
-Finally, we provide a strong yet accessible baseline: a multi-branch convolutional encoder–decoder architecture trained per each task that ingests heterogeneous inputs and predicts task-specific targets.
+Finally, we provide a strong yet accessible baseline: a multi-branch convolutional encoder–decoder architecture inspired by the previous works (Seo et al., 2023, 2024) and largely expanded to ingest heterogeneous inputs and predict our task-specific targets.
+The architecture is trained independently for each task.
 The following list summarizes our contributions:
-  • Benchmark design. We define 14 tasks organized into four groups, with a hierarchical evaluation protocol, standardized windowing and error metrics.
-  • Data packaging. We freeze a stable subset of FAIR-MAST, resolve schema inconsistencies, standardize metadata and units, and release it for reproducibility and long time compatibility.
-  • Tools and API. We provide a Python package for task-specific data loading, processing and batching, masking and alignment utilities, and evaluation logic, integrated with PyTorch stack.
-  • Baseline models.
-We release a multi-branch convolutional encoder–decoder baseline with reference configurations and training scripts, establishing reproducible baselines across all tasks (Seo et al., 2023) (Seo et al., 2024).
-2. Preliminaries
-To ground TokaMark benchmark in the reality of complex heterogeneous fusion data, we first introduce a signal taxonomy that organizes these measurements by function and modality, and allows us to provide a consistent vocabulary for describing inputs, targets, and tasks in the benchmark.
-2.1. Data Taxonomy
+  • Benchmark design. We define 14 tasks organized into four groups, with a hierarchical evaluation protocol, standardized windowing, and error metrics.
+  • Data packaging. We freeze a stable subset of FAIR-MAST data, resolve schema inconsistencies, standardize metadata and units, and release it for reproducibility and long-time compatibility.
+  • Tools and API. We provide a Python package for task-specific data loading, processing and batching, masking and alignment utilities, and evaluation logic, integrated with the PyTorch stack.
+  • Baseline models. We release a multi-branch convolutional encoder–decoder baseline with reference configurations and training scripts, establishing reproducible baselines across all tasks.
 Table 1. Signals taxonomy for TokaMark.
 Category Subcategory/Signals Origin Frequency Modality
-Magnetics Flux loops, pickup coils; saddle coils Sensor 5 kHz; 50 kHz Profile
-Kinetics Thomson scattering; interferometer Sensor 0.2 kHz; 4 kHz Profile; time series
+Magnetics Flux loops, pickup coils; saddle coils Diagnostic 5 kHz; 50 kHz Profile
+Kinetics Thomson scattering; interferometer Diagnostic 0.2 kHz; 4 kHz Profile; time series
 […]
-MHD Precursors Mirnov coils Sensor 500 kHz Profile
-Currents Poloidal Field Coil currents; Solenoid current, Plasma current Sensor 4 kHz Profile; Time series
+Fast magnetics Mirnov coils Diagnostic 500 kHz Profile
+Currents Poloidal field coil currents; solenoid current, plasma current Diagnostic 4 kHz Profile; time series
 Voltages Poloidal field coil voltages Actuator 4 kHz Profile
@@
 Fueling NBI power, gas puffing Actuator 4 kHz Time series
-Equilibrium Shape parameters, $\beta$ metrics; plasma boundary; flux map Derived 0.2 kHz Time series; profile; video
-MAST is a spherically-shaped tokamak located in Culham, Oxfordshire (UK) and was operated by UKAEA and EURATOM from 1999 to 2013 (Sykes et al., 2001; Counsell et al., 2005; Meyer et al., 2009).
-MAST and tokamaks in general operate in short experimental cycles known as discharges or shots.
-The length of those cycles depends on the device size and in the case of MAST they are typically lasting around 2–3 seconds.
-Over its operational lifetime, MAST produced more than 30,000 shots, with each shot containing diagnostic signals measuring various properties of the plasma, including the magnetic field, plasma temperature, shape parameters, applied heating etc.
+Equilibrium Shape parameters, $J_{\text{tor}}$ metrics; plasma boundary; flux map Derived 0.2 kHz Time series; profile; video
+2. Preliminaries
+To ground TokaMark benchmark in the reality of complex heterogeneous fusion data, we first introduce a signal taxonomy that organizes these measurements by function and modality, which allows us to provide a consistent vocabulary for describing inputs, targets, and tasks in the benchmark.
+2.1. Data Taxonomy
+MAST is a spherical tokamak located in Culham, Oxfordshire (UK) and was operated by UKAEA and EURATOM from 1999 to 2013 (Sykes et al., 2001; Counsell et al., 2005; Meyer et al., 2009).
+MAST, and tokamaks in general, operates in short experimental cycles known as discharges or shots.
+The length of those cycles depends on the device size, and in the case of MAST typically last around 2–3 seconds.
+Over its operational lifetime, MAST produced more than 30,000 shots, with each shot containing diagnostic signals measuring various properties of the plasma, including the magnetic field, plasma temperature, shape parameters, and applied heating, to name a few.
 Recent works (Jackson et al., 2025, 2024) have created an open dataset of diagnostic data from a subset of the history of MAST.
-The FAIR-MAST dataset contains 11,573 shots from the last five experimental campaigns on MAST.
-In this work we utilized a total of 39 signals from FAIR-MAST in the design of our benchmark tasks.
-We organize all signals along several complementary axes, summarized in Table 1.
-First, we group signals by category, which encodes their physical semantics–that is, the physical quantities they represent and how they are used by MAST workflows.
+FAIR-MAST data contains 11,573 shots from the last five experimental campaigns on MAST.
+In this work, we utilized a total of 39 signals from FAIR-MAST data for the design of our benchmark tasks.
+We organize all the selected signals along several complementary axes, as summarized in Table 1.
+First, we group signals by category, which encodes their physical semantics—that is, the physical quantities they represent and how they are used by MAST workflows.
 This information corresponds to the first two columns of Table 1.
@@
 For example, the shape parameters subcategory includes several attributes jointly describing plasma geometry.
-A full list of the signals used in this work is provided in the appendix in Table 4.
-Second, we distinguish signals by their origin.
-We define three classes: (i) sensors, corresponding to direct hardware measurements; (ii) actuators, representing controllable machine parameters used to steer plasma behavior; and (iii) derived signals, comprising quantities produced by reconstruction pipelines such as EFIT (e.g., shape parameters or flux maps).
-The third axis is frequency, reflecting the wide range of sampling rates across diagnostics, and spanning from 0.2 kHz up to 500 kHz.
+A full list of the signals used in this work is provided in the appendix in Table A.1.
+Second, we distinguish signals by their origin, for which we define three classes: (i) diagnostics, corresponding to direct hardware measurements; (ii) actuators, representing controllable machine parameters used to steer plasma behavior; and (iii) derived signals, comprising quantities produced by reconstruction pipelines such as EFIT (Appel and Lupelli, 2018) (e.g., shape parameters or flux maps).
+The third axis in Table 1 is frequency, reflecting the wide range of sampling rates across diagnostics, spanning from 0.2 kHz up to 500 kHz.
 Finally, the fourth axis is modality, describing the structural form of each signal: (i) time series, represented as 1D tensors (scalar value over time); (ii) profiles, represented as 2D tensors (vector over time); and (iii) videos, represented as 3D tensors (image over time).
 These latter two axes are particularly important for guiding the design of AI model architectures and loss functions.
-2.2. Tasks Structural Taxonomy
+2.2. Structural Taxonomy of Tasks
 The downstream tasks in TokaMark share a common structural formulation designed to reflect the online, window-based nature of plasma control and forecasting in tokamak experiments.
 Rather than operating on full-shot signals (as is often done in offline post-discharge analysis), our tasks are defined using input window and output window anchored at a reference time point.
-Each task consists of one or more input signals and one or more output signals also referred to as targets.
-Inputs typically include both sensor signals (e.g., magnetics, radiatives, kinetics) and actuator signals (e.g., voltages, fueling), while outputs include either sensor signals or derived quantities like equilibrium reconstructions.
+Each task consists of one or more input signals, and one or more output signals also referred to as targets.
+Inputs typically include both diagnostic signals (e.g., magnetics, radiatives, kinetics) and actuator signals (e.g., voltages, fueling), while outputs include either diagnostic signals or derived quantities like equilibrium reconstructions.
 The alignment between the input and output windows naturally induces several families of modeling objectives:
-  • Reconstruction. Given a set of sensors signals $A$ over the interval $t_{0}-\Delta_{input},t_{0}]$, the goal is to reconstruct a related set of signals $B$ over the same interval.
+  • Reconstruction. Given a set of diagnostic signals $A$ over the interval $t_{0}-\Delta_{input},t_{0}]$, the goal is to reconstruct a related set of signals $B$ over the same interval.
   • Autoregressive (AR) Forecasting.
-Given sensors signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ together with actuator trajectories over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, predict future values of the same signal set over $t_{0},t_{0}+\Delta_{output}]$.
+Given diagnostic signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ together with actuator trajectories over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, predict future values of the same diagnostic signals $A$ over $t_{0},t_{0}+\Delta_{output}]$.
   • Reconstructive (RC) Forecasting.
-Using sensors signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ and actuators over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, forecast a set of related outputs $B$ over $t_{0},t_{0}+\Delta_{output}]$ where $B$ may contain signals as in $A$.
-We further distinguish tasks by their temporal dependency structure: Markovian tasks require only a short input window to make forecasts, reflecting fast dynamics, whereas Non-Markovian (NM) tasks require substantially longer input histories.
-The latter arise in phenomena governed by slower transport processes or evolving stability conditions, where even short-term predictions depend on extended plasma context.
-Finally, TokaMark emphasizes self-supervised and weakly supervised task formulations whenever possible, reducing reliance on expert-labeled targets.
-This design was taken to encourage the development of AI models capable of discovering transferable plasma representations.
+Using diagnostic signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ and actuators over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, forecast a set of related outputs $B$ over $t_{0},t_{0}+\Delta_{output}]$ where $B$ may contain signals as in $A$.
+We further distinguish tasks by their temporal dependency structure: Markovian tasks require only a short input window to make forecasts, reflecting fast dynamics, whereas Non-Markovian (NM) tasks require substantially longer input histories even for short-term predictions.
 2.3. Data-driven Challenges
-To complete the description of the FAIR-MAST dataset we highlight three core data-centric challenges that AI systems must overcome in order to resolve tasks in TokaMark.
+To complete the description of FAIR-MAST data, we highlight three core data-centric challenges that AI systems must overcome in order to resolve tasks in TokaMark.
 These challenges arise directly from the heterogeneous, multi-instrument nature of tokamak diagnostics and the realities of operating large fusion experiments.
 Multi-fidelity.
-Diagnostic systems operate at widely varying sampling rates, ranging from a few hundred hertz to several hundred kilohertz.
+Diagnostic systems operate at varying sampling rates, ranging from a few hundred hertz to several hundred kilohertz.
 Resampling to a common high-frequency time base is computationally expensive and often unnecessary, while down-sampling to the lowest frequency discards information critical for resolving fast plasma phenomena.
@@
 Multi-modality.
-Signals differ not only in physical meaning, units, numerical ranges, but also in structural forms of the corresponding data tensors.
-The dimensionality of the latter can be from 1 to 3 with the number of channel in each non-temporal dimension vary from 1 up to 170.
-AI systems must be able to fuse these diverse modalities into coherent representations suitable for downstream tasks.
+Signals differ not only in physical meaning, units, and numerical ranges, but also in structural forms of the corresponding data tensors.
+Their dimensionality can be from 1 to 3, with the number of channel in each non-temporal dimension varying from 1 up to 170.
+AI systems must be able to fuse these modalities into coherent representations suitable for downstream tasks.
 Missing data.
@@
 Individual signals may also contain missing time segments, for example due to limited acquisition windows or diagnostic failures.
-Naïvely discarding shots or windows with missing components wastes valuable examples and can introduce distributional bias.
-Robust approaches must therefore handle incomplete signals and irregular temporal coverage in a principled manner.
-3. MAST Benchmark
-In this section, we provide a description of the main components of TokaMark: tasks, data preparation and evaluation protocol.
+Naively discarding shots or windows with missing components wastes valuable examples and can introduce distributional bias.
+Robust approaches must therefore handle incomplete signals and irregular temporal coverage.
+3. TokaMark: A MAST Benchmark
+In this section, we provide a description of the main components of TokaMark: tasks, data preparation, and evaluation protocol.
 3.1. Downstream Tasks
 Table 2. Summary of groups and tasks for TokaMark.
-Task Input Sensors Input Actuators Outputs Type Input Window Output Window
-Group 1 1-1 Magnetics, Currents Shape parameters, $\beta$ metrics Reconstruction 5 ms 5 ms
+Task Input Diagnostics Input Actuators Outputs Type Input Window Output Window
+Group 1 1-1 Magnetics, Currents Shape parameters, $J_{\text{tor}}$ metrics Reconstruction 5 ms 5 ms
 1-2 Same as 1-1 Plasma boundary Reconstruction 5 ms 5 ms
 1-3 Same as 1-1 Flux map Reconstruction 5 ms 5 ms
-Group 2 2-1 Magnetics, Currents Voltages, NBI Power Currents, $\beta$ metrics, Shape parameters RC Forecasting 5 ms 25 ms
+Group 2 2-1 Magnetics, Currents Voltages, NBI Power Currents, $J_{\text{tor}}$ metrics, Shape parameters RC Forecasting 5 ms 25 ms
 2-2 Same as 2-1 Same as 2-1 Plasma boundary RC Forecasting 5 ms 25 ms
@@
 3-2 Thomson scattering, Radiatives Same as 3-1 Radiatives AR Forecasting 5 ms 50 ms
-3-3 Magnetics, Currents, Interferometer Same as 3-1 Thomson scattering, $\beta$ metrics RC Forecasting, NM history 5 ms
+3-3 Magnetics, Currents, Interferometer Same as 3-1 Thomson scattering, $J_{\text{tor}}$ metrics RC Forecasting, NM history 5 ms
 Group 4 4-1 Magnetics, Mirnov coils, Currents, Radiatives, interferometer References, Fueling Soft X-ray AR Forecasting, NM history 100 ms
@@
 4-5 Same as 4-1 Same as 4-1 Mirnov diagnostics AR Forecasting, NM history 100 ms
-For the TokaMark we constructed a suite of 14 tasks designed to evaluate the core capabilities expected of AI models for real-world fusion experiments: learning unified representations from heterogeneous diagnostics, temporal reasoning across both fast and slow dynamics, robustness to incomplete state information, and generalisation across operating regimes.
-Those tasks span across diverse timescales, modalities, and predictive objectives.
-We arranged the tasks into 4 broad groups that represent some of the major modeling challenges an AI system must master: instantaneous reconstruction, short‑term magnetics dynamics, slow transport‑driven profile evolution, and long‑range forecasting of rare but safety‑critical events and MHD activity.
+In TokaMark, we assembled a set of 14 tasks spanning across diverse timescales, modalities, and predictive objectives.
+We arranged them into 4 broad groups that represent some of the major modeling challenges arising in real-world fusion experiments.
+The 4 groups include: instantaneous reconstruction, short‑term magnetics dynamics, slow transport‑driven profile evolution, and long‑range forecasting of rare but safety‑critical events and MHD activity.
 Table 2 summarizes definitions of tasks and groups.
 3.1.1. Group 1: Equilibrium Reconstruction
-The tasks in group 1 define a suite of reconstruction problems where the objective is to infer plasma equilibrium – its shape, boundaries, and various properties – from instantaneous magnetic measurements.
-The targets of the three tasks in this group span multiple modalities from scalar parameters (task_1-1) to full plasma contour (task_1-2) and two-dimensional representations of the poloidal magnetic flux (task_1-3).
-The inputs of the tasks are identical and use standard magnetic diagnostics and actuator currents.
+The tasks in Group 1 define a suite of reconstruction problems where the objective is to infer plasma equilibrium—its shape, boundaries, and various properties—from instantaneous magnetic measurements.
+The targets of the three tasks in this group span multiple modalities from scalar parameters (Task 1-1) to full plasma contour (Task 1-2) and two-dimensional representations of the poloidal magnetic flux (Task 1-3).
+The inputs of the tasks are identical and use magnetic diagnostics and actuator coil currents.
 Equilibrium reconstruction is a foundational operation in tokamak experiments, performed routinely after every plasma discharge and forming the basis of essentially all downstream analysis.
-Traditionally, equilibrium is reconstructed by solving an inverse boundary-value problem governed by the Grad–Shafranov equation that describes the balance of magnetohydrodynamic force in axisymmetric tokamak plasmas.
-These solvers require multiple iterations for convergence and rely on parameterized assumptions about plasma profiles, making it challenging to deploy them in real time.
-Group 1 tasks evaluate whether AI models can infer equilibrium structure directly from raw diagnostics, providing fast, numerics-free surrogates for a core analysis task that could underpin both offline interpretation and real-time control applications.
+Traditionally, equilibrium is reconstructed by solving an inverse boundary-value problem governed by the Grad–Shafranov equation that describes the balance of magnetohydrodynamic force in axisymmetric plasmas.
+These solvers, like EFIT++ (Appel and Lupelli, 2018), require multiple iterations for convergence and rely on parameterized assumptions about plasma profiles, making it challenging to deploy them in real time.
+Group 1 tasks evaluate whether AI models can infer equilibrium structure directly from raw diagnostics, providing fast, numerics-free surrogates that could underpin both offline interpretation and real-time control applications.
 3.1.2. Group 2: Magnetics Dynamics
@@
 Conceptually, these tasks parallel those in Group 1, but shift from static reconstruction to sequence-to-sequence forecasting with actuators signals now playing an essential role in the tasks inputs.
-Complexity increases in the three tasks in the group: from forecasting scalar quantities - such as plasma current evolution - to joint prediction of 2D equilibrium geometry over short horizons.
+Complexity increases in the three tasks in the group: from forecasting scalar quantities—such as plasma current evolution—to joint prediction of 2D equilibrium geometry over short horizons.
 At these timescales, plasma evolution is dominated by inductive coupling between active coils, passive conducting structures, and the plasma current, together with the local response of the equilibrium to perturbations.
-These dynamics govern how the plasma reacts to control actions during a discharge and are central to plasma position, shape, and current control.
-While these effects are routinely modeled offline, a model capable of achieving real time latency would enable improved control policies.
-Group 2 probes whether models can learn the effective plasma– machine response encoded in coupled circuit dynamics and plasma response.
+These dynamics govern how the plasma reacts to magnetic control actions during a discharge and are central to plasma position, shape, and current control.
+While these effects are routinely modeled offline, a model capable of achieving real-time latency would enable improved control policies.
+Group 2 probes whether models can learn an effective description of the plasma response to the coupled magnetic dynamics of the active coils and vessel.
 This would directly target capabilities required for closed-loop control, scenario planning, and digital twin applications.
 3.1.3. Group 3: Profile Dynamics
-Group 3 tasks focus on modeling the temporal evolution of kinetic profiles - primarily electron density and temperature - and on forecasting diagnostic signals associated with confinement-mode transitions.
-The three tasks in this group include short-horizon forecasting (task_3-1 and task_3-2) and pseudo-real-time reconstruction using partial, multi-rate diagnostic inputs (task_3-3).
+Group 3 tasks focus on modeling the temporal evolution of kinetic profiles—primarily, electron density and temperature—and on forecasting diagnostic signals associated with confinement-mode transitions.
+The three tasks in this group include short-horizon forecasting (Task 3-1 and Task 3-2) and pseudo-real-time reconstruction using partial, multi-rate diagnostic inputs (Task 3-3).
 Profile evolution is governed by transport physics acting on particle and energy balance, introducing slower characteristic timescales and intrinsic memory effects compared to magnetic dynamics.
-These processes control energy confinement and overall plasma performance and are central to understanding transport, optimising heating schemes, and predicting access to high-performance regimes.
+These processes control energy confinement and overall plasma performance and are central to understanding transport, optimizing heating schemes, and predicting access to high-performance regimes.
 In practice, profile measurements are often sparse, delayed, or unavailable in real time.
-The main goal of Group 3 tasks is to assess AI models ability to integrate incomplete diagnostic information over time to infer latent plasma state variables.
+The main goal of Group 3 tasks is to assess AI models’ ability to integrate incomplete diagnostic information over time to infer latent plasma state variables.
 This would support applications ranging from post-shot analysis, to integrated scenario design and real-time performance monitoring and control.
 3.1.4. Group 4: MHD activity
-Finally, the tasks in Group 4 comprise long-horizon forecasting of thermal quenches (task_4-1 and task_4-2), vertical displacement events (task_4-3), current quenches (task_4-4) and MHD activity leading to Locked Modes (task_4-5).
+Finally, the tasks in Group 4 comprise long-horizon forecasting of thermal quenches (Task 4-1 and Task 4-2), vertical displacement events (Task 4-3), current quenches (Task 4-4), and MHD activity leading to Locked Modes (Task 4-5).
 These tasks are focused on early detection of plasma instability and disruptions precursors that emerge across multiple diagnostic modalities.
-These tasks require processing of rich, high frequency inputs–including magnetics, radiative diagnostics, and time–frequency MHD signatures—over extended temporal windows, exceeding well beyond immediate system response times.
+These tasks require processing of rich, high-frequency inputs—including magnetics, radiative diagnostics, and time–frequency MHD signatures—over extended temporal windows.
 The onset of MHD instabilities and disruptions is tightly linked to the evolving equilibrium and current distribution, even when the nonlinear dynamics of the instability itself are difficult to model explicitly.
-Disruption avoidance and mitigation are critical operational requirements for present-day tokamaks and future reactors, directly affecting machine protection.
+Disruption avoidance and mitigation are critical operational requirements for present-day tokamaks and future reactors, directly affecting machine lifespan.
 In practice, early-warning systems rely on detecting subtle precursors distributed across multiple diagnostics and evolving over extended time windows.
-Group 4 tasks probe models abilities to integrate long-range temporal context and multi-modal information to anticipate loss-of-control events, testing capabilities that are essential for safe and reliable fusion operation.
+Successful performance in Group 4 tasks would demonstrate the ability to integrate long-range temporal context and multi-modal information, allowing the models to anticipate loss-of-control events.
+This capabilities are essential for safe and reliable fusion operation.
+3.1.5. Input and output window length
+For fast magnetic and actuator dynamics (Group 2, tasks 3-1 and 3-2), the observed diagnostics provide a sufficient description of the system state, such that short input windows are often adequate to predict near-term evolution.
+In contrast, profile evolution (task 3-3), confinement transitions, and MHD activity (Group 4) depend on latent plasma state variables—such as current and pressure profiles—that are only partially and indirectly observed.
+For these processes, accurate prediction requires integrating information over extended time intervals, making long input windows essential.
+By explicitly varying the temporal context required across tasks, the benchmark probes a model’s ability to learn both short-timescale system dynamics and long-range temporal dependencies arising from unobserved physics.
 3.2. Data Preparation
-From the FAIR-MAST dataset we take all 11,573 available shots and extract the data for 39 signals required by out task definitions.
-Before being used for model training and evaluation, these signals undergo through a standardized preprocessing pipeline to ensure consistent formatting and alignment.
+We take all 11,573 available shots from FAIR-MAST data and extract the data for 39 signals required by our task definitions.
+Before being used for model training and evaluation, these signals undergo a standardized preprocessing pipeline to ensure consistent formatting and alignment.
 Data split.
 The benchmark dataset is divided into disjoint training, validation, and test subsets.
-To avoid information leakage across sets, the split is performed at the shot level representing independent experiments and employs random sampling technique.
-The split ratio is 80% / 10% / 10%, and all hyperparameter tuning is conducted exclusively on the training and validation sets.
+To avoid information leakage across sets, the split is performed at the shot level representing independent experiments, and employs random sampling technique.
+The split ratio used is 80%/10%/10%, and all hyperparameter tuning is conducted exclusively on the training and validation sets.
 The test set is held out and used only for the final performance evaluation.
 Window segmentation.
-Shot-level signals data is segmented into multiple input and output windows.
-The lengths of those windows are task dependent and specified in Table 2.
+Shot-level signals data is segmented into input and output windows.
+The lengths of those windows are task-dependent as specified in Table 2.
 Windows are extracted using a sliding-window approach with a stride of $0.001$ seconds between consecutive windows.
@@
 We note that signals are used at their original sampling rates, and no resampling or imputation is performed.
-Handling missing information.
+Handling of missing information.
 We impose additional constraints on the test set to ensure meaningful evaluation.
@@
 3.3. Benchmark Evaluation
-As part of the TokaMark, we introduce a hierarchical evaluation protocol for assessing model performance in a manner that aligns with scientific objectives and the structure of FAIR-MAST data.
-The design explicitly separates three levels of the hierarchy: (i) signals (individual physical quantities), and (ii) tasks (well-defined scientific goals) , and (iii) groups (broader physical objectives).
-This ensures that the benchmark assess both low-level prediction quality and high-level scientific utility.
-For low-level evaluation, the hierarchy provides granular signal-level insights that help diagnose which physics regimes the model captures or fails to capture.
-For models trained jointly across multiple tasks, the same hierarchy provides modular evaluations that highlight generality and transfer properties across tasks.
+As part of TokaMark, we introduce a hierarchical evaluation protocol for assessing model performance which explicitly separates three levels of the hierarchy: (i) signals (individual physical quantities), (ii) tasks (well-defined scientific goals), and (iii) groups (broader physical objectives).
+This provides both, signal-level insights that help diagnose which physics regimes the model captures and higher-level scientific utility assessment.
 To respect the hierarchy, the evaluation aggregates errors according to the following progression:
@@
 In the following, $y_{k,i,j}$ denotes the ground truth value and $\hat{y}_{k,i,j}$ the corresponding model prediction.
-Windows are containers of the equal per signal size storing $N_{k}$ samples. We compute a window-level Root-Mean-Square error error as:
+Windows are containers of the equal per-signal size storing $N_{k}$ samples. We compute a window-level Root-Mean-Square error error as:
 (1) $\mathrm{RMSE}_{k,i}=\sqrt{\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}\left(y_{k,i,j}-\hat{y}_{k,i,j}\right)^{2}}$
-Signals serve as first level of the evaluation hierarchy.
-For each signal $k$ containing $M_{k}$ windows, we aggregate all window errors via the signal-level Root-Mean-Square error and normalize it with the empirical standard deviation $\sigma_{k}$ computed for signal $k$ across all evaluation shots:
-(2) $\mathrm{RMSE}_{k}=\sqrt{\frac{1}{M_{k}}\sum_{i=1}^{M_{k}}\mathrm{RMSE}_{k,i}^{2}},\quad\mathrm{NRMSE}_{k}=\frac{\mathrm{RMSE}_{k}}{\sigma_{k}}$
-This yields a dimensionless quantity that expresses prediction error relative to the natural variability of the target making the metric comparable across signals.
+For each signal $k$ for a given task and shot and containing $M_{k}$ windows, we aggregate all window errors via the single-shot signal Root-Mean-Square error and normalize it with the empirical standard deviation $\sigma_{k}$ computed for signal $k$ across all evaluation shots:
+(2) $\mathrm{RMSE}_{k}=\sqrt{\frac{1}{M_{k}}\sum_{i=1}^{M_{k}}\mathrm{RMSE}_{k,i}^{2}},\quad\mathrm{NRMSE}_{k}=\frac{\mathrm{RMSE}_{k}}{\sigma_{k}},$
+which yields a dimensionless quantity that expresses prediction error relative to the natural variability of the target, making the metric comparable across signals.
 An error of $\mathrm{NRMSE}_{k}=1$ corresponds to a model no better than approximating the signal by its mean, while $\mathrm{NRMSE}_{k}<1$ indicates predictive value.
-We also note that $\mathrm{RMSE}_{k}$ and $\mathrm{NRMSE}_{k}$ are computed for a given task and shot.
-Tasks.
-Task-level errors $\tilde{e}_{t}$ combine $K_{t}$ normalized output signal errors into a single score using a uniform mean, allowing one to quantify model performance on the scientific objective as a whole:
+Concerning tasks, task-level errors $\tilde{e}_{t}$ combine $K_{t}$ normalized output signal errors into a single score using a uniform mean, allowing one to quantify model performance on the scientific objective as a whole:
 (3) $\tilde{e}_{t}=\frac{1}{K_{t}}\sum_{k=1}^{K_{t}}\mathrm{NRMSE}_{k}.$
-Shots are the final step of the aggregation sequence.
-They represent distinct plasma discharges and therefore provide natural boundaries for independent experimental realizations.
-We report per-signal and per-task shot-level averaged scores correspondingly defined as
-(4) $\mathrm{SignalScore}=\frac{1}{S}\sum_{s=1}^{S}\mathrm{NRMSE}_{s,k},\quad\mathrm{TaskScore}=\frac{1}{S}\sum_{s=1}^{S}\tilde{e}_{s,t}$
-Here, $S$ is the total number of shots in the test set used for evaluation. Finally, the group level score is taken as an average of the corresponding task scores.
+Shots represent independent experimental realizations, and therefore provide the final step of aggregation.
+We report signal-level and task-level errors as NRMSE based aggregated across shots quantities correspondingly defined as:
+(4) $\text{Signal}_{\mathrm{NRMSE}}=\frac{1}{S}\sum_{s=1}^{S}\mathrm{NRMSE}_{k(s)},\quad\text{Task}_{\mathrm{NRMSE}}=\frac{1}{S}\sum_{s=1}^{S}\tilde{e}_{t(s)}.$
+Here, $S$ is the total number of shots in the test set used for evaluation.
+Finally, the $\text{Group}_{\mathrm{NRMSE}}$ score is taken as an average of the corresponding $\text{Task}_{\mathrm{NRMSE}}$ scores.
 4. Baseline Model Experiments
@@
 Encoders.
-For each input variable $v_{i}\in(L_{t}^{(0,v_{i})}\times L_{h}^{(0,v_{i})}\times L_{w}^{(0,v_{i})})$, an independent convolutional encoder is instantiated according to its modality: 1D convolutions for time series, 2D convolutions for profiles, and 3D convolutions for videos signals.
+For each input variable $v_{i}\in(L_{t}^{(0,v_{i})}\times L_{h}^{(0,v_{i})}\times L_{w}^{(0,v_{i})})$, an independent convolutional encoder is instantiated according to its modality: 1D convolutions for time series, 2D convolutions for profiles, and 3D convolutions for video signals.
 Each encoder consists of a stack of $N$ convolutional layers with kernel size $K$, stride $s$, and padding $p$, followed by ReLU activations, max-pooling, and batch normalization.
@@
 The size of the feature map $L_{d}^{(\ell,v_{i})}$ along dimension $d\in(t,h,w)$ after encoder layer $\ell\in[1,N]$ and the flattened latent representation size $L^{(v_{i})}$ for variable $v_{i}$ can be computed as follows:
-(5) $L_{d}^{(\ell,v_{i})}$ $=\left\lfloor\frac{1}{2}\left\lfloor\frac{L_{d}^{(\ell-1,v_{i})}+2p-K}{s}\right\rfloor+p-\frac{1}{2}\right\rfloor+1$
-(6) $L^{(v_{i})}$ $=2^{N-1}DL_{t}^{(N,v_{i})}L_{h}^{(N,v_{i})}L_{w}^{(N,v_{i})}$
-Figure 2. Multi-branch convolutional encoder–decoder
+(5) $L_{d}^{(\ell,v_{i})}$ $=\left\lfloor\frac{1}{2}\left\lfloor\frac{L_{d}^{(\ell-1,v_{i})}+2p-K}{s}\right\rfloor+p-\frac{1}{2}\right\rfloor+1,$
+(6) $L^{(v_{i})}$ $=2^{N-1}DL_{t}^{(N,v_{i})}L_{h}^{(N,v_{i})}L_{w}^{(N,v_{i})}.$
+Figure 2. Multi-branch convolutional encoder–decoder.
 Latent fusion backbone.
 The latent embeddings produced by all encoder branches are concatenated and passed through a shared backbone.
-This backbone consists of a sequence of linear layers each followed by ReLU activation and dropout, progressively reducing the dimensionality back to $D$.
-This shared representation provides a compact summary of all input data and connects the encoders and decoders blocks.
+This backbone consists of a sequence of linear layers, each followed by ReLU activation and dropout, progressively reducing the dimensionality back to $D$.
+This shared representation provides a compact summary of all input data and connects the encoder and decoder blocks.
 Decoders.
@@
 The sizes $L_{d}^{(\ell,v_{o})}$ of the feature map before decoder layer $\ell\in[0,N-1]$ and $L^{(v_{o})}$ of the compressed flattened feature map are:
-(7) $L_{d}^{(\ell,v_{o})}$ $=\left\lceil s\cdot(L_{d}^{(\ell+1,v_{o})}-1)-2p+K+o_{p}\right\rceil$
-(8) $L^{(v_{o})}$ $=2^{N-1}DL_{t}^{(0,v_{o})}L_{h}^{(0,v_{o})}L_{w}^{(0,v_{o})}$
+(7) $L_{d}^{(\ell,v_{o})}$ $=\left\lceil s\cdot(L_{d}^{(\ell+1,v_{o})}-1)-2p+K+o_{p}\right\rceil,$
+(8) $L^{(v_{o})}$ $=2^{N-1}DL_{t}^{(0,v_{o})}L_{h}^{(0,v_{o})}L_{w}^{(0,v_{o})}.$
 4.2. Experimental Settings
@@
 For our architecture, we adopt a latent embedding dimension of $D=16$, with $N=3$ convolutional layers per encoder and decoder blocks.
-Each convolution uses a kernel size $K=3$, stride $s=3$, padding $p=1$ and output padding $o_{p}=1$.
-The same architectural hyperparameters — number of layers, kernel size, stride, and padding — are applied consistently across all tasks, establishing a unified baseline for the benchmark.
+Each convolution uses a kernel size $K=3$, stride $s=3$, padding $p=1$, and output padding $o_{p}=1$.
+The same architectural hyperparameters—number of layers, kernel size, stride, and padding—are applied consistently across all tasks.
 Data Preprocessing.
-We run a set of model specific preprocessing steps: first, we replace NaN values with zeros, second, we standardize data using signal-computed zero mean and unit variance scaler.
-This ensures numerical stability and helps to improve convergence during training.
-The training and validation data is down-sampled by taking window samples with a stride of $0.005$ seconds for Markovian tasks and $0.025$ seconds for non-Markovian tasks.
-Furthermore, because our architecture requires a fixed-length context, all input windows are truncated to a duration of 50 ms for non-Markovian tasks.
+We run a set of model-specific preprocessing: first, we replace NaN values with zeros, and then we standardize data using signal-computed zero mean and unit variance scaler.
+This ensures numerical stability and helps to improve training convergence.
+The training and validation data is downsampled by taking samples with a stride of $0.005$ seconds for Markovian tasks, and $0.025$ seconds for non-Markovian tasks.
+Furthermore, because our architecture requires a fixed-length context, inputs for non-Markovian tasks are truncated to a duration of $50$ ms.
 Training Procedure.
-We train our model on mini-batches containing data from 32 shots using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and early stopping with a patience of 10 epochs which terminates training if the validation loss does not improve over this period to prevent overfitting.
+We train our model on mini-batches containing data from 32 shots using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and early stopping with a patience of 10 epochs.
 We also employ a multi-output mean squared error loss, which averages the loss across all outputs of the model.
-Table 3 summarizes the size of models per task, computed from the encoder and decoder feature maps.
+Table 3 summarizes models sizes per each task.
 4.3. Experimental Results
-The results of the baseline models evaluation on TokaMark tasks are presented in Table 3 containing task level scores and group averages.
-Those results demonstrate a clear distinction between tasks complexities: the model performs well for equilibrium reconstruction (Group 1) and magnetics dynamics (Group 2) tasks with group-level scores below 0.17 but degrades for profile dynamics (Group 3) and MHD activites (Group 4) tasks confirming that, as expected, these are more difficult tasks.
-Notably, even within the same group tasks scores vary substantially.
-For instance, plasma boundary reconstruction and forecasting tasks 1-2 and 2-2 are resolved much better then their counterparts within the respective groups.
+The results of the baseline model evaluation on TokaMark tasks are presented in Table 3.
+Those results demonstrate a clear distinction between tasks complexities: the model performs well for equilibrium reconstruction (Group 1) and magnetics dynamics (Group 2) tasks with the group-level scores below 0.17, but degrades for profile dynamics (Group 3) and MHD activities (Group 4) tasks, confirming these are more difficult tasks.
+Notably, even within the same group, tasks scores vary substantially.
+For instance, plasma boundary reconstruction and forecasting tasks Task 1-2 and Task 2-2 are resolved much better than their counterparts within the respective groups.
 On the other side, the error score for Task 4-5 is by far the worst and exceeds unity, suggesting the corresponding signals are poorly constrained or inadequately represented.
+Finally, we provide signal level error metrics in the appendix in Table A.2.
 These results should be interpreted with a consideration to the nature of a baseline model: the architecture is generic, without physics-informed priors or task-specific tuning.
-The scores reported – especially those for profile dynamics and MHD activities tasks – highlight intrinsic benchmark difficulty rather than flaws in optimization.
-This baseline establishes a realistic lower bound and identifies areas for improvement, such as physics-aware modeling, task-adaptive weighting, or specialized architectures that should be considered by the users of TokaMark.
-Table 3. TaskScore and GroupScore for TokaMark.
-Task TaskScore Model size (M)
+The scores reported—especially those for profile dynamics and MHD activity tasks—highlight intrinsic benchmark difficulty rather than flaws in optimization.
+Nevertheless, this baseline establishes a realistic lower bound and identifies areas for improvement.
+Table 3. Task and Group errors for TokaMark  tasks.
+$\textbf{Task/Group}_{\mathrm{\textbf{NRMSE}}}$ Parameters(M)
 Task 1-1 0.1882 0.343
@@
 Task 1-3 0.2552 0.257
-Group 1 Score 0.1656
+Group 1 0.163 -
 Task 2-1 0.1551 0.436
@@
 Task 2-3 0.1724 0.295
-Group 2 Score 0.1290
+Group 2 0.1264 -
 Task 3-1 0.3252 0.158
@@
 Task 3-3 0.3607 0.336
-Group 3 Score 0.3380
+Group 3 0.3389 -
 Task 4-1 0.3445 1.518
@@
 Task 4-5 1.0053 4.741
-Group 4 Score 0.4217
+Group 4 0.4761 -
 5. Conclusions
@@
 We believe the adoption of TokaMark will accelerate progress toward practical, data-driven fusion models, foster stronger collaboration between the fusion and machine learning communities, and ultimately contribute to the development of stable and commercially viable fusion energy.
+Appendix A Full Set of Signals
+Table A.1 presents the complete set of signals used in TokaMark.
+The table is aligned with the taxonomy in Table 1 and specifies signals within each category and subcategory.
+We note, that both the data used by TokaMark as well as the FAIR-MAST data are stored in per-shot zarr files.
+Within each zarr file, signals are organized into groups (which are similar, though not identical, to the categories listed in Table A.1).
+We therefore provide each signal’s full identifier in the format ”group name”-”signal name”.
+Appendix B Signal Level Errors
+In addition to the task- and group-level errors we also report more granular signal-level errors in Table A.2.
+Moreover, that table includes not only NRMSE-based signal errors but also RMSE-, MAE- and NMAE-based errors, all computed analogously.
+Table A.1. Full set of signals used in TokaMark.
+Category Subcategory/Signals Full signal name
+Magnetics Flux loops magnetics-flux_loop_flux
+Pickup coils magnetics-[b_field_pol_probe_ccbv_field, b_field_pol_probe_obr_field, b_field_pol_probe_obv_field]
+Saddle coils magnetics-b_field_tor_probe_saddle_voltage
+Kinetics Thomson scattering thomson_scattering-[t_e, n_e]
+Interferometer interferometer-n_e_line
+[…]
+soft X-ray soft_x_rays-[horizontal_cam_lower, horizontal_cam_upper]
+Fast magnetics Mirnov coils magnetics-[b_field_tor_probe_cc_field, b_field_pol_probe_omv_voltage]
+Currents Poloidal field coil currents pf_active-[coil_current, solenoid_current]
+Plasma current summary-ip
+Voltages Poloidal field coil voltages pf_active-coil_voltage
+References Reference plasma current pulse_schedule-i_plasma
+Reference plasma density pulse_schedule-n_e_line
+Fueling NBI power summary-power_nbi
+Gas puffing gas_injection-total_injected
+Equilibrium Shape parameters equilibrium-[elongation, elongation_axis, triangularity_upper, triangularity_lower, x_point_r, x_point_z, minor_radius, magnetic_axis_r, magnetic_axis_z]
+$J_{\text{tor}}$ metrics equilibrium-[q95, beta_tor, beta_pol, beta_normal, bvac_rmag, bphi_rmag]
+plasma boundary equilibrium-[lcfs_r, lcfs_z]
+flux map equilibrium-psi
+Table A.2. Signal errors across all tasks and groups.
+Task Full signal name $\textbf{Signal}_{\mathrm{\textbf{RMSE}}}$ $\textbf{Signal}_{\mathrm{\textbf{MAE}}}$ $\textbf{Signal}_{\mathrm{\textbf{NRMSE}}}$ $\textbf{Signal}_{\mathrm{\textbf{NMAE}}}$
+Group 1 1-1 equilibrium-beta_normal 0.3686 0.2664 0.2734 0.1976
+equilibrium-beta_pol 0.0870 0.0609 0.2668 0.1868
+equilibrium-beta_tor 0.8385 0.6512 0.2415 0.1876
+equilibrium-bphi_rmag 0.0293 0.0251 0.1497 0.1281
+equilibrium-bvac_rmag 0.0230 0.0213 0.1367 0.1270
+equilibrium-elongation 0.0501 0.0414 0.2342 0.1936
+equilibrium-elongation_axis 0.0564 0.0472 0.2676 0.2240
+equilibrium-magnetic_axis_r 0.0147 0.0117 0.1557 0.1237
+equilibrium-magnetic_axis_z 0.0105 0.0086 0.1429 0.1173
+equilibrium-minor_radius 0.0114 0.0091 0.1805 0.1437
+equilibrium-q95 0.5950 0.4729 0.1749 0.1390
+equilibrium-triangularity_lower 0.0223 0.0181 0.2277 0.1851
+equilibrium-triangularity_upper 0.0220 0.0178 0.2155 0.1744
+equilibrium-x_point_r 0.1099 0.0799 0.1431 0.1041
+equilibrium-x_point_z 0.0926 0.0599 0.0130 0.0084
+1-2 equilibrium-lcfs_r 0.0211 0.0147 0.0559 0.0390
+equilibrium-lcfs_z 0.0262 0.0197 0.0404 0.0303
+1-3 equilibrium-psi 0.0142 0.0050 0.2552 0.0901
+Group 2 2-1 equilibrium-beta_normal 0.3102 0.2210 0.2301 0.1639
+equilibrium-beta_pol 0.0734 0.0509 0.2252 0.1562
+equilibrium-beta_tor 0.7493 0.5718 0.2158 0.1647
+equilibrium-bphi_rmag 0.0225 0.0178 0.1152 0.0911
+equilibrium-bvac_rmag 0.0181 0.0163 0.1079 0.0969
+equilibrium-elongation 0.0460 0.0373 0.2147 0.1742
+equilibrium-elongation_axis 0.0533 0.0441 0.2527 0.2090
+equilibrium-magnetic_axis_r 0.0138 0.0109 0.1462 0.1155
+equilibrium-magnetic_axis_z 0.0102 0.0079 0.1393 0.1080
+equilibrium-minor_radius 0.0106 0.0084 0.1677 0.1324
+equilibrium-q95 0.5151 0.4151 0.1514 0.1220
+equilibrium-triangularity_lower 0.0214 0.0170 0.2187 0.1734
+equilibrium-triangularity_upper 0.0214 0.0168 0.2093 0.1641
+equilibrium-x_point_r 0.0976 0.0656 0.1272 0.0854
+equilibrium-x_point_z 0.0870 0.0557 0.0122 0.0078
+pf_active-coil_current 535.13 374.30 0.1057 0.0739
+pf_active-solenoid_current 913.93 749.73 0.0726 0.0595
+summary-ip 26035.74 21395.01 0.0803 0.0660
+2-2 equilibrium-lcfs_r 0.0212 0.0149 0.0561 0.0393
+equilibrium-lcfs_z 0.0307 0.0230 0.0473 0.0355
+2-3 equilibrium-psi 0.0096 0.0037 0.1724 0.0663
+Group 3 3-1 thomson_scattering-n_e $4.73\times 10^{18}$ $3.43\times 10^{18}$ 0.3307 0.2399
+thomson_scattering-t_e 107.00 71.96 0.3196 0.2149
+3-2 soft_x_rays-horizontal_cam_lower 0.00719 0.00343 0.2382 0.1136
+soft_x_rays-horizontal_cam_upper 0.00735 0.00355 0.2717 0.1313
+spectrometer_visible-filter_spectrometer_dalpha_voltage 0.2172 0.1226 0.4830 0.2726
+3-3 equilibrium-beta_normal 0.4389 0.2981 0.3256 0.2211
+equilibrium-beta_pol 0.0987 0.0628 0.3028 0.1926
+equilibrium-beta_tor 0.9948 0.7296 0.2865 0.2101
+thomson_scattering-n_e $7.35\times 10^{18}$ $5.69\times 10^{18}$ 0.5143 0.3984
+thomson_scattering-t_e 125.35 94.33 0.3744 0.2817
+Group 4 4-1 soft_x_rays-horizontal_cam_lower 0.00962 0.00523 0.3184 0.1731
+soft_x_rays-horizontal_cam_upper 0.01003 0.00562 0.3705 0.2077
+4-2 soft_x_rays-horizontal_cam_lower 0.00971 0.00488 0.3215 0.1616
+soft_x_rays-horizontal_cam_upper 0.00922 0.00485 0.3406 0.1792
+4-3 equilibrium-magnetic_axis_z 0.0198 0.0171 0.2702 0.2341
+4-4 summary-ip 139237.10 91351.51 0.4292 0.2816
+4-5 magnetics-b_field_pol_probe_omv_voltage 0.1595 0.0683 0.9371 0.4014
+magnetics-b_field_tor_probe_cc_field $2.15\times 10^{-5}$ $5.95\times 10^{-6}$ 1.0734 0.2972